sqrt T + 12 - T = 0

T+12-T=0
sqrt(T+12)-T=0
sqrt(T+12)=X
(sqrtT+12)^2=T^2
T+12=T^2
T^2-T-12=0

I came up with (T = 4)(T-3) and I don't if I'm right please help!

Answer by oscargut(686)

if equation is:
sqrt (T + 12) - T = 0 (eq 1)
then you did it well and
T^2-T-12=0 (eq 2)
T=4 and T=-3
but solutions of the (eq 1) are included in solutions of (eq 2)
(-3 is a solution of the eq 2 but not a solution of eq 1)
Answer: T=4